# sample.bib

### v1

@@ 4.2 Generative Methods
Min et al.[min2023uniworld] 2023 Image-LiDAR pairs 4D GO UL/SSL This work proposes a spatial-temporal World Model for unified autonomous driving pretraining.
@@
VP Finn et al.[173] 2016 Videos Videos UL/SSL This work develops an action-conditioned video prediction model for interaction that incorporates appearance information in previous frames with motion predicted by the model.
Micheli et al.[174] 2022 Images, Actions Images SSL The World Model is composed of a discrete autoencoder and an autoregressive transformer. And it is trained in a self-supervised manner.
@@
$\bullet$ This “AD” refers to “Autonomous Driving”and this “VP” refers to “Visual Prediction”. $\bullet$ This “GO” refers to “Geometric Occupancy” and this “OG” refers to “Occupancy Grids”. $\bullet$ This “SSL” refers to “Self-Supervised Learning” and this “UL” refers to “Unsupervised Learning”.
@@ 4.3 Non-Generative Methods
In contrast to generative World Models, LeCun[135] elaborated his different conception of World Model by proposing a Joint Extraction and Prediction Architecture (JEPA) based on Energy-based Model. This is a non-generative and self-supervised architecture, as it does not predict the output $y$ directly from the input $x$, but encodes $x$ as $sx$ to predict $sy$ in an abstract space. This has the advantage that it does not have to predict all the information about $y$, and can eliminate irrelevant details.
@@
Since its proposal, the JEPA architecture has been applied by several scholars in different fields with excellent performance. In computer vision, Assran et al.[176] proposed I-JEPA for learning highly semantic image representations without relying on manual data enhancement. Combining I-JEPA with ViT resulted in strong downstream performance in a variety of tasks ranging from linear classification to object counting and depth prediction. Bardes et al. proposed MC-JEPA in Rrf.[177], which uses the JEPA architecture and a self-supervised learning approach to achieve co-learning of optical flow and content features from video to learning dynamic content features. MC-JEPA performs well in various tasks such as optical flow estimation, and segmentation of images and videos. In the graph domain, Skenderi et al. proposed Graph-JEPA[178], which is a JEPA model for graph domains. It divides the input graph into subgraphs and then predicts the representation of the target subgraph in the context subgraph. Graph-JEPA has obtained excellent performance in both graph classification and regression problems.
JEPA is a self-supervised learning architecture based on Energy-based Model that observes and learns the laws of how the world works and is highly generalizable. There is also great potential for JEPA in autonomous driving, which is expected to generate high-quality driving scenarios and driving strategies by learning how driving works.
@@ 5 Conclusion and Future Directions
This paper provides a comprehensive overview of FMs in autonomous driving. Recent works on the application of FMs, notably language and vision models, towards advancing autonomous driving are summarized in detail in Section 2, and their involvement in shaping the research of end-to-end autonomous driving is extensively surveyed. Recent works on data augmentation of FMs are detailed in Section 3. In Section 4 we present an exploratory application of the World Model in autonomous driving. Overall, FMs can effectively contribute to autonomous driving in terms of both data enhancement and model optimization.
@@
However, many of the aforementioned studies also raised concerns about the maturity of FMs. Their suitability for the safety-critical downstream task of autonomous driving requires further development and is awaiting confirmation. Whether it is the hallucinations[zhang2023siren, liu2023exposing], security concerns such as inference latency[pope2023efficiently, weng2023inference], limited context length[liu2023lost, longchat2023], or practical issues such as storage limitations, deployment costs, and ethical and legal constraints that have yet to be established, the solution of “autonomous driving through foundation models” to achieve autonomous driving on public roads across various weather and traffic conditions still requires a comprehensive and holistic approach. Among them, the hallucination error problem is mainly manifested as misrecognition in autonomous driving, such as wrong target detection, which may cause serious safety accidents. The phantom problem mainly arises because of the limited samples in the dataset or because the model is affected by unbalanced or noisy data, and the stability and generalization ability of the model needs to be strengthened by means of expanding data and adversarial training. Challenges in security are also severe, while inference latency may result in vehicles not being able to make timely decisions and execute them, posing serious traffic safety risks. FMs have seen an amazing emergent ability with model scaling, and demonstrated success in natural language processing. However, in the context of autonomous driving, this line of research faces additional open challenges due to limited available data and extended context length issues. These challenges contribute to an inadequate understanding of macroscopic driving scenarios, thereby complicating long-term planning in this field. Driving video is a high-dimensional continuous modal with an extremely large amount of data (several orders of magnitude larger compared to textual data). Hence, training large visual model requires a more macroscopic scene distribution to embed enough video frames to reason about complex dynamic scenes, which requires a more robust network structure and training strategies to learn this information. In a recent study, Bai et al.[bai2023sequential] proposed a two-stage approach, in which images are converted into discrete tokens to obtain “visual sentences”, and then autoregressive predictions are made, similar to the standard approach for language models[11]. Another promising solution may lie in World Models. As described in Section 4, World Models can learn the intrinsic evolutionary laws of the world by observing a small number of events that are either relevant or irrelevant to the task. However, World Models also have certain limitations in exploratory applications, where uncertainty in the predictive outcomes of the models, as well as learning what kind of data captures the intrinsic laws of how the world works, still warrant further exploration. Multi-modal data fusion is expected to provide more feature information and may be a promising research direction.
From the findings of related research, another direction worth further exploration is how to design an effective architecture for self-supervised learning. Self-supervised learning can effectively break through the limitation of data labeling, allowing models to learn data at scale and fully unleash the inference capability of the model. If a Foundation Model for autonomous driving can be trained using scaled driving scenario data under the paradigm of self-supervised learning, it is anticipated that its generalization and emergent capability will be greatly improved. This advancement may lead to the realization of a more generalized Foundation Model. In conclusion, although there are many challenges to be solved in applying Foundation Models to autonomous driving, its potential has already begun to take shape. In future, we will continue to monitor the progress of Foundation Models applied to autonomous driving.

### v2

@@ 4.2 Generative Methods
Zheng et al.[zheng2023occworld] 2023 3D Occupancy Scene Scene, Ego-vehicle Motion SSL By constructing a 3D occupancy space, a World Model is trained to predict the next scene from previous Scenarios, following an autoregressive manner. This work utilizes VQ-VAE for discretizing the 3D occupancy scene into tokens.
AD Min et al.[min2023uniworld] 2023 Image-LiDAR pairs 4D GO UL/SSL This work proposes a spatial-temporal World Model for unified autonomous driving pretraining.
@@
VP Finn et al.[173] 2016 Videos Videos UL/SSL This work proposed to interact with the world under unsupervised conditions and develops an action-conditioned model for video prediction.
@@
Wang et al.[wang2024worlddreamer] 2024 Images, Videos, Text, Actions Videos SSL Visual inputs are mapped into discrete tokens using VQ-GAN, and then the masked tokens are predicted using Transformer.
$\bullet$ This “AD” refers to “Autonomous Driving” and this “VP” refers to “Visual Prediction”. $\bullet$ This “GO” refers to “Geometric Occupancy” and this “OG” refers to “Occupancy Grids”.
@@ 4.3 Non-Generative Methods
In contrast to generative World Models, LeCun[135] elaborated on different conceptions of World Model by proposing a Joint Extraction and Prediction Architecture (JEPA) based on energy-based model. This is a non-generative and self-supervised architecture, as it does not predict the output $y$ directly from the input $x$, but encodes $x$ as $sx$ to predict $sy$ in representation space, as illustrated in Fig.8. This has the advantage that it does not have to predict all the information about $y$ and can eliminate irrelevant details.
Figure 8: Comparison of the architecture of generative and non-generative methods[176].
Since its proposal, the JEPA architecture has been applied by several scholars in different domains with excellent performance. In the graph domain, Skenderi et al.[178] proposed Graph-JEPA, which is a JEPA model for graph domains. It divides the input graph into subgraphs and then predicts the representation of the target subgraph in the context subgraph. Graph-JEPA has obtained excellent performance in both graph classification and regression problems. In the field of audio, Huang et al.[fei2023jepa] proposed A-JEPA, which applies the mask modeling principle to audio. Following experimental validation, A-JEPA has been demonstrated to perform well in speech and audio classification tasks. Sun et al. proposed JEP-KD[sun2024jep], which employs an advanced knowledge distillation method to enhance the effectiveness of Visual Speech Recognition (VSR) and narrow the performance gap between VSR and Automatic Speech Recognition (ASR).
@@
In the field of computer vision, Bardes et al.[177] proposed MC-JEPA, which employs the JEPA architecture and a self-supervised learning approach to facilitate co-learning of optical flow and content features, thereby enabling the acquisition of dynamic content features. From video, MC-JEPA performs well in a variety of tasks, including estimation of optical flow, and segmentation of images and videos. META[176] proposed I-JEPA for learning highly semantic image representations without relying on manual data enhancement. The combination of I-JEPA with Vision Transformers yielded strong downstream performance in a variety of tasks, including linear classification, object counting, and depth prediction. Building on I-JEPA, META applies JEPA to the video domain by proposing V-JEPA[bardes2024revisiting]. This method combines mask prediction with the JEPA architecture to train a series of V-JEPA models with feature prediction as the goal of self-supervised learning. Experimental results demonstrate that these models exhibit excellent performance in a range of computer vision downstream tasks, including action recognition, action classification, and target classification.
To date, no literature has been identified that directly applies the JEPA to the field of autonomous driving. Nevertheless, it has great potential. Firstly, instead of predicting the video in pixel space, non-generative world models make feature predictions in representation space. This eliminates many irrelevant details. For example, in the scene prediction task of autonomous driving, we are more interested in the future movements of other traffic participants on the current road. Furthermore, for other vehicles that are not on the current road of the autonomous vehicle, for example, situated next to an elevated road parallel to the current road, we do not consider their future motion trajectories. The JEPA model eliminates these irrelevant details and reduces the complexity of the problem. Additionally, V-JEPA has demonstrated its ability to learn features in video. By analyzing a sufficiently large number of driving videos, it is anticipated that V-JEPA will be widely used in tasks such as generating driving scenarios and predicting future environmental states.
5 Data Augmentation Based on Foundation Models
As deep learning continues to evolve, the performance of FMs with pre-training and fine-tuning as the underlying architecture is improving. FMs are spearheading the transition from rule-driven to data-driven learning paradigms. The significance of data as a key aspect of model learning is evident. A substantial quantity of data is utilized in the training process of an autonomous driving model to facilitate the model’s comprehension and decision-making abilities in diverse driving scenarios. Nevertheless, the collection of realistic data is a time-consuming and laborious process, so data augmentation is crucial to improving the generalization ability of automatic driving models.
The realization of data augmentation needs to consider two aspects: on the one hand, how to obtain large-scale data so that the data fed to the autonomous driving system is diverse and extensive; on the other hand, how to obtain as much high-quality data as possible so that the data which is used to train and test the autonomous driving models is accurate and reliable. Related works have also roughly chosen two directions to enhance the autonomous driving data, one is to enrich the data content of existing datasets and enhance the data features of driving scenarios, and the other is to generate driving scenarios with multiple levels through simulation. In the following, a review of related works on enhancing data based on FMs will be presented, in Section 5.1 we describe related work on extending datasets, and in Section 5.2 we describe related work on generating driving scenarios. Table LABEL:tab:5 provides a brief overview of some representative works.
5.1 Expansion of Autonomous Driving Datasets
Existing autonomous driving datasets are mostly obtained by recording sensor data and then labeling the data. The features of the data obtained in this are usually low-level and exist more at the level of numerical representation, which is insufficient for the visuo-spatial features of the autonomous driving scenarios. Natural language descriptions are seen as an effective way to enhance scene representation[4], Flickr30k[108], RefCOCO[109], RefCOCOg[110], and CLEVR-Ref[111] use concise natural language descriptions to identify the corresponding visual regions in an image. Talk2Car[112] fused image, radar, and LiDAR data to construct the first object-referenced dataset containing language commands for self-driving cars. However, the Talk2Car dataset allowed only one object to be referenced at a time. CityFlow-NL[113] constructed a dataset for multi-target tracking through natural language descriptions, and Refer-KITTI[114] achieved prediction of arbitrary target tracking by leveraging natural language queries in the corresponding task.
FMs provide new ideas for enriching and expanding autonomous driving datasets under their advanced semantic understanding, reasoning, and interpretation capabilities. Qian et al.[115] created NuScenes-QA, a visual question-and-answer dataset for autonomous driving in 3D multi-view driving scenarios, by encoding question descriptions through a language model and obtaining answers through feature fusion with sensor data. Significant progress was made in the use of natural language prompts. Wu et al.[116] extended NuScenes-QA by constructing the dataset NuPrompt by capturing and combining natural language elements and then invoking LLM to generate the descriptions. The dataset provided a finer match between the 3D instances and each of the prompts, which helped to characterize objects in the autopilot images more accurately. Sima et al.[drivelm] took into account the interactions of the traffic elements and constructed Graph Visual Question Answering by extending the nuScenes dataset[128] with BLIP-2, which can better clarify the logical dependencies between objects and the hierarchy of driving tasks. In addition to directly extending the augmented autonomous dataset, some scholars have also integrated the Chain-of-Thought (CoT) capability of LLM and the cross-modal capability of the vision model to build an automatic annotation system, OpenAnnotate3D[117], which can be used for multi-modal 3D data. Expanding the dataset by utilizing the advanced understanding, reasoning, and interpretation capabilities of the underlying models can help to better assess the interpretability and control of the autonomous driving system, thus improving the safety and reliability of the autonomous driving system. A comparison of some representative work is shown in Table 3.
Table 3: Comparison of extended datasets. ‘-’ means unavailable.
Dataset Source Based FMs Modality 3D Multi-Views VIdeos Frames QA Pairs
RefCOCO[110] COCO None image referring expression × × - 26711 -
Refer-KITTI[114] KITTI None image point cloud object referral × × 18 6650 -
Talk2Car[112] nuScences None image point cloud driving command × - 9217 -
nuPrompt[116] nuScences GPT-3.5 image point cloud Question-Answering 850 34149 35k
DriveLM-nuScences[drivelm] nuScences BLIP-2 image point cloud Question-Answering - 4871 443k
5.2 Generation of Driving Scenarios
The diversity of driving scenarios is of great significance for autonomous driving. To obtain a better generalization ability, autonomous driving models must learn a wide variety of scenarios. However, the reality is that driving scenarios conform to a long-tailed distribution (It is a probability distribution in which a significant proportion of the observations or instances are concentrated in the tail(s) of the distribution, away from the center or mean.). The “long-tail problem” of autonomous driving vehicles is that they are capable of handling situations that are frequently encountered, but are unable to cope with corner cases in rare or extreme situations. To address the long-tail problem, the key is to get as many corner cases as possible. Nevertheless, it is inefficient to limit the collection to real scenarios. For instance, in CODA[li2022coda], work on corner case mining, there are only 1,057 valid data out of 1 million data.
Given the above, the generation of large-scale and high-quality driving scenario data necessitates the capacity to actively create a multitude of driving scenarios. Traditional methodologies may be classified into two primary categories: rule-based and data-driven. Rule-based approaches, as exemplified by the literature cited in references[118, 119, 120, 121], necessitate the utilization of predefined rules, are inadequate for the characterization of complex environments, simulate simpler environments, and exhibit limited generalization ability. In contrast, data-driven approaches[122, 123, 124, 125] utilize driving data to train the model, enabling it to continuously learn and adapt. However, data-driven approaches often necessitate a substantial quantity of labeled data for training, impeding further development of driving scenario generation. Additionally, this approach lacks control and is unsuitable for custom generation. Recently, FMs have achieved considerable success, and the generation of higher-quality driving scenarios through FMs has also attracted significant research attention. On the one hand, the diversity and accuracy of data generation can be enhanced based on the powerful understanding and reasoning capabilities of FMs. On the other hand, diverse prompts can be designed for controlled generation.
Table 4: Video generation performance on nuScenes dataset. ‘-’ means unavailable. The FID indicator and FVD indicator provide feedback on the image and video quality, respectively.
Method Based FMs Multi-View Multi-Frame FID↓ FVD↓
BEVGen[swerdlow2024street] None × 25.54 -
DriveGAN[kim2021drivegan] None 73.4 502.3
MagicDrive[gao2024magicdrive] CLIP × 16.20 -
Panancea[wen2023panacea] CLIP 16.96 139
DriveDreamer[164] WM 52.6 452.0
DriverDreamer-2[zhao2024drivedreamer] GPT-3.5 & WM 11.2 55.7
Driving-WM[wang2023driving] WM 15.8 122.7
Based on LLMs and VLMs. In response to the fact that some long-tailed scenarios can never be collected in multi-view shots, Yang et al.[130] fused verbal cues, BEV sketch, and multi-view noise to design a two-stage generative network BEVControl for synthesizing realistic street scene images. Nevertheless, BEVControl is insufficient for modeling foreground and background detail information. To address the difficulty of obtaining large-scale BEV representations, Li et al.[131] proposed a spatio-temporal consistent diffusion framework, DrivingDiffusion, to autoregressively generate realistic multi-view videos controlled by 3D layouts. The quality of the generated data can be effectively enhanced by introducing local cue inputs into the vision model. For controllable generation, Wen et al.[wen2023panacea] integrated text prompts, image conditions, and BEV sequences to design a controllable module to improve the controllability of driving scenarios generation. Gao et al.[gao2024magicdrive] designed 3D geometric control by integrating text prompts with camera pose, road map, and object box fusion control to generate diverse road scenarios.
Based on the powerful understanding and reasoning ability of LLMs and VLMs, it has also become a research hotspot to embed them directly or guide the model to generate driving scenarios. Marathe et al.[129] efficiently generated a dataset comprising 16 weather extremes via prompting leveraging a VLM. Nevertheless, the model had some extension limitations due to the phenomenon of pre-selected fixation in data selection. Chen et al.[132] realized the combination of numerical vector modality and natural language by pairing control commands collected by reinforcement learning intelligence and question answers generated by LLM to directly construct new data. Zhong et al.[133] proposed a scenario-level diffusion-based language-guided traffic simulation model, CTG++, which can generate instruction-compliant, realistic, and controllable traffic scenarios. Wang et al.[105] utilized natural language descriptions as conceptual representations that were integrated with LLM to enrich the complexity of the generated scenarios by leveraging their powerful common-sense reasoning capabilities. The behavior of human drivers is also an important part of driving scenarios, Jin et al.[134] proposed SurrealDriver, a generative driving agent simulation framework for urban environments based on LLM. By analyzing and learning from real driving data, SurrealDriver can capture the driver’s behavior patterns and decision-making processes and generate behavior sequences that are similar to those in real driving.
Based on World Models. To achieve the controllable generation of driving scenarios, Wang et al.[164] combine text prompts and structured traffic constraints to guide the generation of pixel points with text descriptions. To obtain more accurate dynamic information, Wang et al.[wang2023driving] incorporate driving actions into a controllable architecture, utilizing text descriptors, layouts, and ego actions to control video generation. However, these approaches introduce more structural information, which limits the interactivity of the model. To address this issue, Zhao et al.[zhao2024drivedreamer] propose a novel approach that combines LLM with World Model. This approach involves using LLM to convert user queries into agents’ trajectories, which are then used to generate HDMap. This HDMap then guides the generation of driving videos.
Efficient and accurate controllability generation can be achieved using FMs for driving scenarios. This will be able to provide diverse training data, which is important for improving the generalization ability of autonomous driving systems. A comparison of some representative work is shown in Table 4. Furthermore, the generated driving scenarios can be used to evaluate different autonomous driving models to test and validate their performance. Of course, we should also be able to see that with the emergence of various large-scale FMs such as Sora and Genia, there are new potential ideas for the generation of autonomous driving videos. The models are not restricted to the driving domain but can be used for transfer learning utilizing models obtained from training in the general vision domain. While the current state of technology in this domain remains imperfect, we believe that in the future, with the breakthrough of related technologies, we can even use them to generate the various driving scenarios we need, and truly learn a World Model to simulate the world.
Table 5: Works on Data Augmentation
Authors Input Output Learning
Expand Dataset Qian et al.[115] 2023 Images, Text, Point Clouds Q-A Pairs SL
Wu et al.[116] 2023 Images, Text Object-Text Pairs SL
Zhou et al.[117] 2023 Images, Text Labeled Data SL
Generate Scenarios Marathe et al.[129] 2023 Objects, Scenarios, Weather Condition Multi-Weather Images SL
Wen et al.[wen2023panacea] 2023 Text, BEV Sequence Multi-View Videos SL
Chen et al.[132] 2023 Objects, Text Trajectory SL
Generate Scenarios Zhong et al.[133] 2023 Text, Scenarios with Noise Traffic Scenarios SL
Wang et al.[105] 2023 Images, Knowledge Latent Space Simulation SL
Zhao et al.[zhao2024drivedreamer] 2024 Text Videos SL
$\bullet$ This “Q-A” refers to “Question-Answer”.
@@ 6 Conclusion and Future Directions
This paper provides a comprehensive overview of the application of FMs to autonomous driving. In Section 3, recent works on the application of FMs such as LLMs and VLMs to autonomous driving are summarized in detail. In Section 4, we present an exploratory application of the World Models to the field of autonomous driving. In Section 5, recent works on data augmentation of the FMs are detailed. Overall, the FMs can effectively assist autonomous driving in terms of both augmenting the data and optimizing the model.
To evaluate the effectiveness of FMs in autonomous driving, we compare different FMs with traditional methods in terms of their effectiveness in motion planning in Table 6. Due to the relative maturity of LLMs and VLMs, it can be observed that methods based on them to enhance autonomous driving have been improved overall. In contrast, WMs-based approaches are still undergoing further exploration, with relatively less work published. Nevertheless, through the previous analysis, we can also see that World Models are excellent at learning the evolutionary laws of the physical world and have great potential for improving autonomous driving.
Table 6: Motion planning performance on the nuScenes validation dataset. †: Results of perception and prediction from UniAD. ‡: Results of perception and prediction from VAD. *: Results of perception and prediction from dataset annotations.
Method Based FMs L2(m)↓ Collision(%)↓
1s 2s 3s Avg. 1s 2s 3s Avg.
ST-P3[hu2022st] None 1.33 2.11 2.90 2.11 0.23 0.62 1.27 0.71
UniAD[104] None 0.48 0.96 1.65 1.03 0.05 0.17 0.71 0.31
VAD[jiang2023vad] None 0.41 0.70 1.05 0.72 0.07 0.17 0.41 0.22
GenAD[zheng2024genad] None 0.36 0.83 1.55 0.91 0.06 0.23 1.00 0.43
GPT-Driver[72] LLM 0.21 0.43 0.79 0.48 0.16 0.27 0.63 0.35
GPT-Driver[72]* LLM 0.20 0.42 0.72 0.44 0.14 0.25 0.60 0.33
Agent-Driver[mao2023language] LLM 0.22 0.65 1.34 0.74 0.02 0.13 0.48 0.21
DriveVLM-Dual[tian2024drivevlm] VLM 0.17 0.37 0.63 0.39 0.08 0.18 0.35 0.20
DriveVLM-Dual[tian2024drivevlm] VLM 0.15 0.29 0.48 0.31 0.05 0.08 0.17 0.10
VLP-UniAD[pan2024vlp] LLM 0.36 0.68 1.19 0.74 0.03 0.12 0.32 0.16
VLP-VAD[pan2024vlp] LLM 0.30 0.53 0.84 0.55 0.01 0.07 0.38 0.15
OccWorld-O[zheng2023occworld]* WM 0.43 1.08 1.99 1.17 0.07 0.38 1.35 0.60
Drive-WM[wang2023driving] WM 0.43 0.77 1.20 0.80 0.10 0.21 0.48 0.26
Challenges and Future Directions. Nevertheless, it is evident from previous studies that FM-based autonomous driving is not yet sufficiently mature. This phenomenon can be attributed to several factors. FMs suffer from the problem of hallucination[zhang2023siren, liu2024exposing], as well as the fact that there are still limitations in learning video, a high-dimensional continuous modality. Additionally, deployment issues caused by inference latency[pope2023efficiently, weng2023inference] and potential ethical implications and societal impact also be considered.
Hallucination. The hallucination error problem is mainly manifested as misrecognition in autonomous driving, such as wrong target detection, which may cause serious safety accidents. The phantom problem mainly arises because of the limited samples in the dataset or because the model is affected by unbalanced or noisy data, and the stability and generalization ability needs to be strengthened by utilizing expanding data and adversarial training.
Real-world deployment. As previously discussed, the majority of current research on FMs in autonomous driving is based on open-source dataset experiments[tian2024drivevlm, pan2024vlp] or closed-loop experiments in simulation environments[lmdrive, wang2023drivemlm], which is insufficient for real-time considerations. Additionally, some studies[pope2023efficiently, weng2023inference] have highlighted that large models have certain inference latency, which could potentially lead to significant safety concerns in autonomous driving applications. To further explore the effectiveness of FMs for real-time applications in autonomous driving, we conducted an experimental[ww]. We used LoRA to fine-tune the LLaMA-7B[44], and the fine-tuned LLM can reason to generate driving language commands. To verify its real-time performance in driving scenarios, we reasoned on a single GPU A800 and a single GPU 3080, respectively, and the time required to generate 6 tokens is 0.9s and 1.2s, effectively verifying that vehicle deployment of FM is possible. To further utilize the computing resources in the vehicle, other modules, such as the intelligent cockpit, can be considered to be deployed in the cloud. Therefore, the final form of deployment in the real world can be speculated: deploying the intelligent cockpit module in the cloud, while the intelligent driving module is directly deployed in the vehicle, is a relatively more reasonable choice. In the future, with the improvement of edge computing and in-vehicle computing capabilities[xu2023sparks], it may gradually move to a hybrid deployment model of vehicle, road, and cloud to further improve real-time response capabilities and privacy protection.
AI Alignment. The deepening of FMs into various industries, including autonomous driving, is a significant trend. Nevertheless, as related research continues, so do the risks to human society. Advanced AI systems exhibiting undesirable behaviors (e.g., spoofing) are a cause for concern, especially in areas such as autonomous driving, which is directly related to personal safety, and requires serious discussion and reflection. In response to this, AI Alignment has been proposed and is currently being developed. The objective of AI Alignment is to align the behaviors of AI systems with human intentions and values. This approach focuses on the goals of AI systems rather than their capabilities[leike2018scalable]. AI Alignment facilitates the risk control, operational robustness, human ethicality, and interpretability of advanced AI systems when implemented in various domains[ji2024ai]. This is a substantial body of research encompassing numerous AI-related disciplines. As this paper concentrates on the domain of autonomous driving and does not delve into the specifics of risk causes and solutions, we will not elaborate further here. In the field of autonomous driving, it is important to note that while promoting the application of FMs, researchers must establish reasonable technical ethics based on the guidance of AI Alignment. This includes paying attention to the issues of algorithmic fairness, data privacy, system security, and the human-machine relationship. Furthermore, it is essential to promote the unity of technological development and social values to avoid potential ethical and social risks.
@@
Visual Emergent Abilities. FMs have seen amazing emergent abilities with model scaling and demonstrated success in natural language processing. Nevertheless, in the context of autonomous driving, this line of research faces additional open challenges due to limited available data and extended context length issues. These challenges contribute to an inadequate understanding of macroscopic driving scenarios, thereby complicating long-term planning in this field. Driving video is a high-dimensional continuous modal with an extremely large amount of data (several orders of magnitude larger compared to textual data). Hence, training large vision models requires a more macroscopic scene distribution to embed enough video frames to reason about complex dynamic scenarios, which requires a more robust network structure and training strategies to learn this information. Bai et al.[bai2023sequential] proposed a two-stage approach in a recent study, in which images are converted into discrete tokens to obtain “visual sentences”, and then autoregressive predictions are made, similar to the standard approach for language model[11]. Another promising solution may lie in World Models. As described in Section 4, World Models can learn the intrinsic evolutionary laws of the world by observing a small number of events that are either relevant or irrelevant to the task. However, World Models also have certain limitations in exploratory applications, where uncertainty in the predictive outcomes of the models, as well as learning what kind of data captures the intrinsic laws of how the world works, still warrant further exploration.
In conclusion, although there are many challenges to be solved in applying Foundation Models to autonomous driving, its potential has already begun to take shape. In the future, we will continue to monitor the progress of Foundation Models applied to autonomous driving.